# Superintegrability of the monomial Uglov matrix model

### v1

@@ 4.2 Representation of the content product in terms of Uglov polynomials
As announced in the introduction another result of this paper is the representation of (49) in terms of Uglov polynomials evaluated at a special locus. This is an important part of the superintegrability relation. In other known cases of superintegrability in matrix models the whole answer is expressed through special values of symmetric functions. A similar effect was not known even for the undeformed non-gaussian monomial model. Here we conjecture an answer for $\beta$-deformed contents in terms of Uglov polynomials. It’s analog for Schur polynomials immediately follows by setting $\beta=1$ in the formulas below and is a particular case of known specializations formulas. We believe that the respective formulas for Uglov polynomials can be proven starting directly from the Macdonald polynomials, but we don’t do it in this paper.
The product depends not only on $a$ but also on $N\mod s$. In fact it depends only the difference $N-a\mod s$. Therefore it makes sense to choose a representative with $a=0$. Formulas with other $a$ can be obtained from it. Here we propose the following formula:
@@
In the generic case the corresponding locus in terms of $p_{k}$ variables will be given by multiple root of unity sums, while it looks much simpler in the $x_{i}$ variables. Formulas for $[[\beta N-\beta i+j]]_{s,a}$ can be obtained in a similar fashion. One way to do it is to switch to $p_{k}$ variables, where one can formally substitute $N$ by $N-1+\frac{1}{\beta}$, i.e.:
@@
One can get the corresponding formulas for Schur functions simply by taking the $\beta\rightarrow 1$ limit. Therefore in this section we have demonstrated that the content products of the non-Gaussian model can be represented as Uglov/Schur polynomials are special points restoring a stronger version of superintegrability in this model. On important point to make here is that this statement is actually also in the spirit of the next section. The fact that the simplest $\beta$-deformation of the non-Gaussian content product requires precisely Uglov polynomials is a strong argument in favor of the validity of the model.
@@ 5 In support of the conjecture
In this section we provide various arguments for the validity of our conjecture and the relevance of the root of unity limit for non-Gaussian matrix models. In particular in sec. 5.4 we explain the main reasoning that brought us to consider this conjecture.
@@ 5.1 Explicit evaluation of the integrals
The most straightforward way to try to test or look for superintegrability is to compute expectation values directly at fixed $N$ and $\beta$ and try to come up with the proper symmetric functions. Consider the simplest not trivial example is the one with parameters $r=2,s=2$. First, notice that averages in odd degree vanish. Therefore it makes sense to compute only correlators with even degree. A simple basis to do the calculations initially is the power sums basis. Therefore at level two we get
@@
This expression quite explicitly demonstrates that superintegrability is a rather non-trivial statement - choosing the wrong basis leads to completely intractable formulas. In the case at hand it’s possible to make the correct guess and deduce the level $4$ Uglov polynomials:
@@
It appears that at $v=0$ the model can be reduced to the standard $\beta$-deformation at least when the expectation values are non-zero. Therefore in the non-vanishing sector of the $v=0$ model superintegrability can be proved. As stated above, non-zero expectation values correspond to $v=0$ Uglov polynomials for partitions with all parts divisible by $s$. Consider such a case:
@@
After this change of variables the all the rays are ”glued” together:
@@
The integral is convergent for $Re(\nu)>-1$, which in our case is satisfied. Therefore substituting $\nu=\frac{1-s}{s}\,,a_{1}=\frac{1}{s}$ and $\beta\rightarrow\tilde{\beta},R\rightarrow\tilde{R}$ we can present :
@@
Combining the results above we obtain exactly the desired formula.
@@ 5.3 Virasoro constraints
One of the powerful tools in the theory of matrix models are the Virasoro constraints [58, 59, 53]. They can be presented as a set of relations between correlators or as differential equations for the times extended partition function. In both cases their origin is the independence of the integral under the change of integration variables. Relation of Virasoro constraints to superintegrability was studied in [60]. However, that work was limited to a few simple cases. In the case at hand Virasoro constraints for generic potential were studied in [37]. Already in this case it appears that, contrary to the standard story only a certain subset of constrains can be constructed. On the other hand, it is well known, that when one specifies the potentials, additional features might appear. This specifically happens in the cases of non-Gaussian monomial (and polynomial) potentials. In that case even in the standard Hermitian models, the Virasoro constraints do not completely determine the partition function, but have ambiguities related to contour choices.
@@
Now, in the Gaussian model, these constraints completely fix the form of the partition function as a formal series, and therefore, by formula (82) determine the Schur expectation values. On the other hand already in the cubic model this is not the case and the system of Virasoro constraints has a non trivial kernel given by generic functions of $p_{1}$. This subject is covered in detail in [53].
@@
The derivation for a generic potential, just as in (85), was done in [37]. The key takeaway for us if that due to the structure of the Uglov determinant, only a subset of equations can be obtained. In particular for the constraint generated by $x_{i}\to x_{i}+\epsilon x_{i}^{n+1}$ the variation of the Uglov determinant contains:
@@
The Virasoro constraints for the the monomial Uglov model are obtained by a redefinition of variables similar to (81). However, now, to accommodate for the expansion in terms of Uglov polynomials, it also makes sense to normalize the times in such way that we obtain the correct Cauchy formula (40). That is, we set in the Virasoro constrains (88):
@@
Here $N$ and $\beta$ can be treated simply as parameters and we are not forced to interpolate between discrete values. As a result we are bounded only by complexity of calculating Uglov polynomial, and hence Macdonald polynomials, for bigger representations. We checked our conjecture from $r=2$ to $r=5$ at least for $|R|\leq 10$. Surely one can go even to higher orders.
@@ 5.4 $W$-representation
Another important piece of evidence for the conjecture, and the main source of intuition for us, is the consideration of relevant $W$-operators. In this section we will briefly remind the main construction of $W$-operators for matrix models and then see where it leads when applied to the monomial matrix models. The main idea behind $W$-operators is that they are a sort of hidden symmetry of the matrix model, that is supposedly responsible for superintegrability. This point of view is explored in [30], where it was also explained how to construct $W$-operators for specific models. The starting point of this subject is the observation [61] that times dependent partition functions of some matrix models can be represented in the form:
@@
The operator $\hat{W}$ acts on parameters $p_{k}$ and depends on the model: on the potential, measure and type of matrices integrated. In some cases it is either more appropriate or necessary that the exponent acts not on an identity, but on some other functions like in the WLZZ models [62, 63] where it acts on an exponent with additional parameters. Recently some progress was made in understanding the role of $W$-operators in superintegrability, and it was revealed that one should study their action on the relevant symmetric functions. In this regard one should view these $W$-operators as parts of some large algebra that has distinguished representations in terms of partitions. Such algebras, now studied under the name of BPS-algebras, include the $\mathcal{W}_{1+\infty}$ algebra, affine Yangians $Y(\hat{\mathfrak{gl}}_{n})$ etc. [24, 22, 23, 25, 26, 27, 28]. In regards to matrix models only some of these algebras have yet appeared. As a result of this section we will observe the relevance of supposedly the $Y(\hat{\mathfrak{gl}}_{n})$ Yangian.
It should be noted that this representation can in some cases be derived from Virasoro constraints (or higher $W$-constraints). In these situations the aforementioned action on partitions can be used to prove superintegrability, à la [29]. In other cases it is not possible, or rather more peculiar, for instance, when Virasoro constraints do not completely fix the partition functions. The monomial and Uglov models are exactly of this sort. Nevertheless, classic techniques combined with recent progress on the matter allow to reverse engineer the $W$-representation if superintegrability is already known, which then opens up the possibility to use methods coming from representation theory of relevant algebras. This is exactly the approach we are going to take here.
@@
at $q=\omega_{s}$. At this point we can already observe the root of unity limit emerging, which in fact this is the key observation of this section. Now we can go on to explicitly represent the function (97) as well and then build the $W$-operator. We will not go on to provide a full expression, as we don’t yet have a concise way to present is. However it is clear, that the building block of this $W$-operator will be an operator acting on Schur functions as:
@@
and hence we can explicitly write it’s second quantized bosonic representation:
@@
Which is nothing but the Macdonald operator (normalized) at $q=t$ [64]. Therefore the $W$-operator for specific monomial matrix model with be made out of iterated commutators of the Macdonald operator evaluated at:
@@
At this locus the Macdonald operator always acts on Schur functions, since for any value of $q=t$ Macdonald polynomials reduce to Schur polynomials. Therefore the symmetric function in these models are not sensitive to the specific point in the $(q,t)$-plane. However, we clearly see, that the operators are different and moreover, we could clearly expect that deformation around this point would depend on the specific value of $q$. Following this logic we are lead to search for the $\beta$-deformation around the root of unity locus, which turns out to be the correct thing to do as demonstrated in the previous sections. Finally, let us note that this $\beta$-deformation of the root of unity point has recently been studied from the purely algebraic standpoint, at least, in the case of $r=2$. It was shown that the resulting algebra is the affine Yangian $Y(\hat{\mathfrak{gl}}_{2})$ and the respective analogs of cut-and-join operators were constructed [41, eq.(97)-(100)].
We point out once again, that these consideration where not derived from Virasoro constraints, since in non-Gaussian examples it is not possible, or at least not known at the moment. Hence in this section we described the logic that was announced in fig 1.
@@
Here we would to explore the possibility to derive not only the Uglov determinant, but the whole matrix model together with the potential from the $(q,t)$-deformed integrals. To do this let us first recall that it is conjectured that the $(q,t)$-analogs of the Gaussian and Complex matrix (RCM) models are also exactly solvable. To define the models we first recall some notations.
@@
What this discussion suggests is that one can recover the Gaussian matrix model from the $(q,t)$-deformed WL ensemble (107) in the limit $q\rightarrow\omega_{2}$. On the other hand one take the standard $q\rightarrow 1$ limit of the Gaussian $(q,t)$ model(106) instead. By similar consideration one would have:
@@
Therefore we clearly can observe that we obtain two versions of the Gaussian matrix model. One with an Uglov determinant with $r=s=2$ and the other with $r=1\,,\,s=2$. We have observed in sec. 4 that both case are superintegrable, which is in agreement with the limits.
@@ 6 Discussion and outlook
One thing that we want to stress is that our result fills in an important gap in the list of superintegrable/exactly solvable matrix models. Moreover, from this perspective we obtained a rather interesting example. Suppose we were to look at the case at hand from the perspective of finding the right basis for a given measure and potential. Given the rather complicated form of the Uglov determinant, guessing an answer might seem improbable. Moreover no other methods for solving the model work in this case, so we can’t even calculate a number of averages by some analytic method and then try to combine those to obtain factorized formulas. Hence it is quite surprising that an answer does exist. On the other hand we observe that for a given $s$ only a special potential produces nice formulas, further stressing that superintegrability is a peculiar property.
@@
  • Proving superintegrability algebraically. As we have explained our proposal can’t be proven directly from Virasoro constraints, since there is no obvious way to obtain the $W$-operator from them. This is not merely a technical difficulty but a conceptual one. Virasoro constraints do not fix the integration contour, while superintegrability in the presented form seems to strongly depend on it. On the other hand, Virasoro constraints and $W$-operators are among the few tools for matrix models that survive $\beta$-deformation.
@@
  • There are several technical extensions that are expected based on previous results. First is the case of non-normalized averages à la [56]. Clearly we expect that a distinguished role will be played by Uglov polynomials for rectangular partitions, which interplays nicely with [39]. Second is the extension towards bilinear correlators along the lines of [66].
@@ Appendix A Symmetric functions and notations
Integer partitions and Young diagrams are denoted by $R=[R_{1},R_{2},R_{3},\ldots]$. Superintegrability formulas often are given in terms of the so called contents of the partitions. Is is defined in terms of the coordinates of a box in the Young diagram $(i,j)$ with $i$ being the row number and $j$ the column number. For example the shaded box in the figure below has coordinates $(2,3)$:

### v2

@@ 4.2 Representation of the content product in terms of Uglov polynomials
As announced in the introduction, another result of this paper is the representation of (49) in terms of Uglov polynomials evaluated at a special locus. This is an important part of the superintegrability relation. In other known cases of superintegrability in matrix models, the whole answer is expressed through special values of symmetric functions. A similar effect was not known, even for the undeformed non-gaussian monomial model. Here we conjecture an answer for $\beta$-deformed contents in terms of Uglov polynomials. Its analog for Schur polynomials immediately follows by setting $\beta=1$ in the formulas below and is a particular case of known specializations formulas. We believe that the respective formulas for Uglov polynomials can be proven starting directly from the Macdonald polynomials, but we don’t do it in this paper.
The product depends not only on $a$ but also on $N\mod s$. In fact, it depends only the difference $N-a\mod s$. Therefore, it makes sense to choose a representative with $a=0$. Formulas with other $a$ can be obtained from it. Here we propose the following formula:
@@
In the generic case the corresponding locus in terms of $p_{k}$ variables will be given by multiple root of unity sums, while it looks much simpler in the $x_{i}$ variables. Formulas for $[[\beta N-\beta i+j]]_{s,a}$ can be obtained similarly. One way to do it is to switch to $p_{k}$ variables, where one can formally substitute $N$ by $N-1+\frac{1}{\beta}$, i.e.:
@@
One can get the corresponding formulas for Schur functions simply by taking the $\beta\rightarrow 1$ limit. Therefore, in this section we have demonstrated that the content products of the non-Gaussian model can be represented as Uglov/Schur polynomials are special points restoring a stronger version of superintegrability in this model. One important point to make here is that this statement is actually also in the spirit of the next section. The fact that the simplest $\beta$-deformation of the non-Gaussian content product requires precisely Uglov polynomials is a strong argument in favor of the validity of the model.
@@ 5 In support of the conjecture
In this section, we provide various arguments for the validity of our conjecture and the relevance of the root of unity limit for non-Gaussian matrix models. In particular, in sec. 5.4 we explain the main reasoning that brought us to consider this conjecture.
@@ 5.1 Explicit evaluation of the integrals
The most straightforward way to try to test or look for superintegrability is to compute expectation values directly at fixed $N$ and $\beta$ and try to come up with the proper symmetric functions. Consider the simplest, not trivial example is the one with parameters $r=2,s=2$. First, notice that averages in odd degree vanish. Therefore, it makes sense to compute only correlators with even degree. A simple basis to do the calculations initially is the power sums basis. Therefore at level two we get
@@
This expression quite explicitly demonstrates that superintegrability is a rather non-trivial statement: choosing the wrong basis leads to completely intractable formulas. In the case at hand it’s possible to make the correct guess and deduce the level $4$ Uglov polynomials:
@@
It appears that at $v=0$ the model can be reduced to the standard $\beta$-deformation at least when the expectation values are non-zero. Therefore, in the non-vanishing sector of the $v=0$ model superintegrability can be proved. As stated above, non-zero expectation values correspond to $v=0$ Uglov polynomials for partitions with all parts divisible by $s$. Consider such a case:
@@
After this change of variables all the rays are ”glued” together:
@@
The integral is convergent for $Re(\nu)>-1$, which in our case is satisfied. Therefore, substituting $\nu=\frac{1-s}{s}\,,a_{1}=\frac{1}{s}$ and $\beta\rightarrow\tilde{\beta},R\rightarrow\tilde{R}$ we can present :
@@
Combining the results above, we obtain exactly the desired formula.
@@ 5.3 Virasoro constraints
One of the powerful tools in the theory of matrix models are the Virasoro constraints [58, 59, 53]. They can be presented as a set of relations between correlators or as differential equations for the times extended partition function. In both cases, their origin is the independence of the integral under the change of integration variables. Relation of Virasoro constraints to superintegrability was studied in [60]. However, that work was limited to a few simple cases. In the case at hand Virasoro constraints for generic potential were studied in [37]. Already in this case it appears that, contrary to the standard story, only a certain subset of constrains can be constructed. On the other hand, it is well known, that when one specifies the potentials, additional features might appear. This specifically happens in the cases of non-Gaussian monomial (and polynomial) potentials. In that case even in the standard Hermitian models, the Virasoro constraints do not completely determine the partition function, but have ambiguities related to contour choices.
@@
Now, in the Gaussian model, these constraints completely fix the form of the partition function as a formal series, and therefore, by formula (82) determine the Schur expectation values. On the other hand, already in the cubic model this is not the case and the system of Virasoro constraints has a non-trivial kernel given by generic functions of $p_{1}$. This subject is covered in detail in [53].
@@
The derivation for a generic potential, just as in (85), was done in [37]. The key takeaway for us is that due to the structure of the Uglov determinant, only a subset of equations can be obtained. In particular, for the constraint generated by $x_{i}\to x_{i}+\epsilon x_{i}^{n+1}$ the variation of the Uglov determinant contains:
@@
The Virasoro constraints for the monomial Uglov model are obtained by a redefinition of variables similar to (81). However, now, to accommodate for the expansion in terms of Uglov polynomials, it also makes sense to normalize the times in such a way that we obtain the correct Cauchy formula (40). That is, we set in the Virasoro constrains (88):
@@
Here $N$ and $\beta$ can be treated simply as parameters, and we are not forced to interpolate between discrete values. As a result, we are bounded only by complexity of calculating Uglov polynomial, and hence Macdonald polynomials, for bigger representations. We checked our conjecture from $r=2$ to $r=5$ at least for $|R|\leq 10$. Surely one can go even to higher orders.
@@ 5.4 $W$-representation
Another important piece of evidence for the conjecture, and the main source of intuition for us, is the consideration of relevant $W$-operators. In this section, we will briefly remind the main construction of $W$-operators for matrix models and then see where it leads when applied to the monomial matrix models. The main idea behind $W$-operators is that they are a sort of hidden symmetry of the matrix model, that is supposedly responsible for superintegrability. This point of view is explored in [30], where it was also explained how to construct $W$-operators for specific models. The starting point of this subject is the observation [61] that times dependent partition functions of some matrix models can be represented in the form:
@@
The operator $\hat{W}$ acts on parameters $p_{k}$ and depends on the model: on the potential, measure, and type of matrices integrated. In some cases it is either more appropriate or necessary that the exponent acts not on an identity, but on some other functions like in the WLZZ models [62, 63] where it acts on an exponent with additional parameters. Recently some progress was made in understanding the role of $W$-operators in superintegrability, and it was revealed that one should study their action on the relevant symmetric functions. In this regard one should view these $W$-operators as parts of some large algebra that has distinguished representations in terms of partitions. Such algebras, now studied under the name of BPS-algebras, include the $\mathcal{W}_{1+\infty}$ algebra, affine Yangians $Y(\hat{\mathfrak{gl}}_{n})$ etc. [24, 22, 23, 25, 26, 27, 28]. In regard to matrix models only some of these algebras have yet appeared. As a result of this section we will observe the relevance of supposedly the $Y(\hat{\mathfrak{gl}}_{n})$ Yangian.
It should be noted that this representation can in some cases be derived from Virasoro constraints (or higher $W$-constraints). In these situations the aforementioned action on partitions can be used to prove superintegrability, à la [29]. In other cases it is not possible, or rather more peculiar, for instance, when Virasoro constraints do not completely fix the partition functions. The monomial and Uglov models are exactly of this sort. Nevertheless, classic techniques combined with recent progress on the matter allow reversing engineer the $W$-representation if superintegrability is already known, which then opens up the possibility to use methods coming from representation theory of relevant algebras. This is exactly the approach we are going to take here.
@@
at $q=\omega_{s}$. At this point we can already observe the root of unity limit emerging, which in fact this is the key observation of this section. Now we can go on to explicitly represent the function (97) as well and then build the $W$-operator. We will not go on to provide a full expression, as we don’t yet have a concise way to present is. However, it is clear, that the building block of this $W$-operator will be an operator acting on Schur functions as:
@@
and hence we can explicitly write its second quantized bosonic representation:
@@
Which is nothing but the Macdonald operator (normalized) at $q=t$ [64]. Therefore the $W$-operator for a specific monomial matrix model with be made out of iterated commutators of the Macdonald operator evaluated at:
@@
At this locus the Macdonald operator always acts on Schur functions, since for any value of $q=t$ Macdonald polynomials reduce to Schur polynomials. Therefore, the symmetric function in these models are not sensitive to the specific point in the $(q,t)$-plane. However, we clearly see, that the operators are different and moreover, we could clearly expect that deformation around this point would depend on the specific value of $q$. Following this logic we are led to search for the $\beta$-deformation around the root of unity locus, which turns out to be the correct thing to do as demonstrated in the previous sections. Finally, let us note that this $\beta$-deformation of the root of unity point has recently been studied from the purely algebraic standpoint, at least, in the case of $r=2$. It was shown that the resulting algebra is the affine Yangian $Y(\hat{\mathfrak{gl}}_{2})$ and the respective analogs of cut-and-join operators were constructed [41, eq.(97)-(100)].
We point out once again, that these considerations were not derived from Virasoro constraints, since in non-Gaussian examples it is not possible, or at least not known at the moment. Hence in this section we described the logic that was announced in fig 1.
@@
Here we would to explore the possibility to derive not only the Uglov determinant, but the whole matrix model together with the potential from the $(q,t)$-deformed integrals. To do this, let us first recall that it is conjectured that the $(q,t)$-analogs of the Gaussian and Complex matrix (RCM) models are also exactly solvable. To define the models we first recall some notations.
@@
What this discussion suggests is that one can recover the Gaussian matrix model from the $(q,t)$-deformed WL ensemble (107) in the limit $q\rightarrow\omega_{2}$. On the other hand one take the standard $q\rightarrow 1$ limit of the Gaussian $(q,t)$ model(106) instead. By similar considerations, one would have:
@@
Therefore, we clearly can observe that we obtain two versions of the Gaussian matrix model. One with an Uglov determinant with $r=s=2$ and the other with $r=1\,,\,s=2$. We have observed in sec. 4 that both case are superintegrable, which is in agreement with the limits.
@@ 6 Discussion and outlook
One thing that we want to stress is that our result fills in an important gap in the list of superintegrable/exactly solvable matrix models. Moreover, from this perspective we obtained a rather interesting example. Suppose we were to look at the case at hand from the perspective of finding the right basis for a given measure and potential. Given the rather complicated form of the Uglov determinant, guessing an answer might seem improbable. Moreover, no other methods for solving the model work in this case, so we can’t even calculate a number of averages by some analytic method and then try to combine those to obtain factorized formulas. Hence it is quite surprising that an answer does exist. On the other hand, we observe that for a given $s$ only a special potential produces nice formulas, further stressing that superintegrability is a peculiar property.
@@
  • Proving superintegrability algebraically. As we have explained, our proposal can’t be proven directly from Virasoro constraints, since there is no obvious way to obtain the $W$-operator from them. This is not merely a technical difficulty but a conceptual one. Virasoro constraints do not fix the integration contour, while superintegrability in the presented form seems to strongly depend on it. On the other hand, Virasoro constraints and $W$-operators are among the few tools for matrix models that survive $\beta$-deformation.
@@
  • There are several technical extensions that are expected based on previous results. First is the case of non-normalized averages à la [56]. Clearly, we expect that a distinguished role will be played by Uglov polynomials for rectangular partitions, which interplays nicely with [39]. Second is the extension towards bilinear correlators along the lines of [66].
@@ Appendix A Symmetric functions and notations
Integer partitions and Young diagrams are denoted by $R=[R_{1},R_{2},R_{3},\ldots]$. Superintegrability formulas are often given in terms of the so-called contents of the partitions. It is defined in terms of the coordinates of a box in the Young diagram $(i,j)$ with $i$ being the row number and $j$ the column number. For example, the shaded box in the figure below has coordinates $(2,3)$: